# SReferences

### 25 Mar 2024 v1

@@ Abstract
Figure 3: Dynamics of magnetisation exhibiting DTC, DTQC, and thermalisation. The raw data $\langle\hat{Z}_{\rm avg}(t)\rangle_{0}$ (cyan circles) and the error-mitigated data $\langle\hat{Z}_{\rm avg}(t)\rangle$ (yellow diamonds) obtained on the heavy-hexagonal lattice of $L=133$ qubits at various parameter sets $(\theta_{x},\theta_{z})$: a $(0.9\pi,0)$, b $(0.9\pi,0.5\pi)$, c $(0.9\pi,\pi)$, d $(0.85\pi,0)$, e $(0.85\pi,0.5\pi)$, f $(0.85\pi,\pi)$, g $(0.8\pi,0)$, h $(0.8\pi,0.5\pi)$, i $(0.8\pi,\pi)$, j $(0.7\pi,0.0\pi)$, k $(0.7\pi,0.5\pi)$, and l $(0.7\pi,\pi)$. Notice that the time duration in the horizontal axis in j-l is half of that in the other panels.
In addition to the period-doubling oscillation, the longitudinal field $\theta_{z}$ induces a longer-period oscillation, as clearly seen in Fig. 3 (also see Fig. S6). To analyse this additional oscillation, we perform a discrete Fourier transform of the error-mitigated magnetisation, defined as $\tilde{Z}(\omega)=|\frac{1}{n_{\rm max}}\sum_{n=0}^{n_{\rm max}-1}\langle\hat{Z}_{\rm avg}(nT)\rangle e^{-i\omega nT}|$, where $\omega T=2\pi k/n_{\rm max}$ with $k=0,1,\cdots n_{\rm max}-1$ represents the discrete frequency and the number of time steps involved in the Fourier transform is set to $n_{\rm max}=100$. As shown in Fig. 4a-c, when $\theta_{z}=0$, only a single peak at $\omega T/(2\pi)=0.5$ appears in the Fourier spectrum, indicating the presence of the sole period-doubling DTC. Introducing the longitudinal field $\theta_{z}$ leads to the appearance of additional side peaks symmetrically at $\omega_{\pm}$ on both sides of the main peak. These side peaks gradually deviate from the main peak as $\theta_{z}$ increases. Although $\omega_{\pm}T/(2\pi)$ are always rational numbers of the form $k/n_{\rm max}$ by definition, their systematic dependence on $\theta_{z}$ suggests that each of $\omega_{\pm}T/(2\pi)$ will fluctuate around a generally irrational number for a fixed set of parameters when the number of time steps involved in the Fourier transform is varied. The emergence of these side peaks at frequencies $\omega_{\pm}$, which are susceptible to changes in microscopic parameters and essentially incommensurate with the driving period, serves as a signature of a DTQC [22, 25].
@@
Figure 4: Fourier analysis of DTQC frequencies induced by the longitudinal field. Fourier spectrum $\tilde{Z}(\omega)$ of the averaged magnetisation for a $\theta_{x}=0.9\pi$, b $\theta_{x}=0.85\pi$, and c $\theta_{x}=0.8\pi$ with various values of $\theta_{z}=0,0.125\pi,0.25\pi,\cdots,\pi$ (from bottom to top). The results are offset by 0.05 for adjacent values of $\theta_{z}$ for better visibility. d, Envelope frequency $\omega_{\text{env}}$ of the observed DTQCs as a function of the perturbation parameter $\varepsilon$ to the transverse field. e, Intensity of the side peaks, $A_{\text{side}}$, as a function of $\theta_{z}$. A schematic phase diagram showing the crossover between a DTC and a DTQC is provided at the top of the figure. Error bars in a-c and e represent the propagated error due to sampling errors in $\tilde{Z}(\omega)$. However, they are smaller than the size of symbols.
@@
The envelope frequency of the longer-period DTQC oscillation can be estimated as $\omega_{\text{env}}=(\omega_{+}-\omega_{-})/2$. We observe that the envelope frequency increases proportionally to the perturbation to the transverse field $h_{x}$, i.e., $\omega_{\text{env}}\propto\epsilon$ (Fig. 4d). Particularly at $\theta_{z}=\pi$, $\omega_{\text{env}}T/(2\pi)=0.05$, $0.07$, and $0.09$ for $\theta_{x}=0.9\pi$, $0.85\pi$, and $0.8\pi$, respectively. These values approximately follow $\omega_{\text{env}}T/(2\pi)\simeq\epsilon/(2\theta_{J})$, which is proportional to the frequency of a Bloch oscillation [43] induced by a weak transverse field in a one-dimensional Ising model with long-range interactions [44].
@@
To characterise the crossover between a DTC and a DTQC, we plot the sum of the intensities $A_{\pm}$ of the side peaks at $\omega_{\pm}$, denoted as $A_{\text{side}}=A_{+}+A_{-}$, in Fig. 4e. $A_{\rm side}$ tends to decrease with decreasing $\theta_{x}$. At $\theta_{x}=0.7\pi$, $\langle\hat{Z}_{\rm avg}(t)\rangle$ quickly decays with increasing time step $t/T$ as seen in Figs. 3j-l, indicating the absence of both DTC and DTQC. When $\theta_{x}=\pi$, only the period-doubling DTC is present, as $|\langle\hat{Z}_{\rm avg}(t)\rangle|$ remains 1 regardless of the value of $\theta_{z}$. Therefore, we conclude that DTQCs manifest in a prethermal state persisting within the timescale $0\leqslant t/T\leqslant 100$ for a parameter range of $0.8\pi\lesssim\theta_{x}\lesssim 0.9\pi$ and $0.25\pi\lesssim\theta_{z}\leqslant\pi$.
@@
The recent advancements in the quality of quantum devices have enabled successful simulation of quantum dynamics on a much larger scale, both in terms of qubit count and time step duration. Specifically, achieving quantum dynamics simulation on a 133-qubit system for up to 100 time steps using a digital quantum computer, as demonstrated in this study, represents a significant leap forward compared to previous studies. Our comparative study, employing a 2dTNS method, confirmed the agreement between the results obtained using the quantum device and those from the classical simulations, for approximately up to 50 time steps. However, in our investigation of both 28-qubit and 133-qubit systems, we encountered a critical parameter regime where the perturbation to the transverse field approaches $2\epsilon\approx 0.2\pi$, near the crossover boundary between DTCs/DTQCs and thermalisation. In this regime, the limitation of classical simulations become apparent, particularly for long-time steps around 100. Indeed, in the 28-qubit system, we observe discrepancies between the results of the 2dTNS method, even with the largest feasible bond dimension, and those of the state-vector simulation (Supplementary Information S2). Additionally, in the 133-qubit system, a slow convergence of the results of the 2dTNS method with respect to the bond dimension is observed, implying the growth of entanglement that is hardly tractable within the classical resources available for the 2dTNS method (Supplementary Information S3). These observations highlight the boundary of the classical simulation capabilities. Therefore, our study of Floquet dynamics in this parameter regime, extending up to 100 time steps on the 133-qubit system, pushes the boundaries of classical simulations to their limit, emphasising the significant potential of current digital quantum computers for simulating out-of-equilibrium quantum dynamics in two dimensions.
This work is based on results obtained in part from a project, JPNP20017, subsidised by the New Energy and Industrial Technology Development Organization (NEDO), Japan. We acknowledge the support form the Japan Society for the Promotion of Science (JSPS) KAKENHI Grants (Grant Nos. JP19K23433, JP21H04446, JP22K03479, JP22K03520, and JP23K13066) from the Ministry of Education, Culture, Sports, Science and Technology (MEXT), Japan. We also appreciate the funding received from JST COI-NEXT (Grant No. JPMJPF2221) and the Program for Promoting Research of the Supercomputer Fugaku (Grant No. MXP1020230411) from MEXT, Japan. Furthermore, we acknowledge the support from the UTokyo Quantum Initiative, the RIKEN TRIP project, and the COE research grant in computational science from Hyogo Prefecture and Kobe City through the Foundation for Computational Science. Tensor-network simulations are based on the high-performance tensor computing library GraceQ/tensor [45]. A part of the numerical simulations has been performed using the HOKUSAI supercomputer at RIKEN and the supercomputer Fugaku installed at RIKEN Center for Computational Science.
Supplementary Information to “Unveiling clean two-dimensional discrete time quasicrystals on a digital quantum computer”
@@
${}^{1}$omputational Quantum Matter Research Team, RIKEN Center for Emergent Matter Science (CEMS), Wako, Saitama 351-0198, Japan
${}^{2}$uantum Computational Science Research Team, RIKEN Center for Quantum Computing (RQC), Wako, Saitama 351-0198, Japan
${}^{3}$omputational Materials Science Research Team, RIKEN Center for Computational Science (R-CCS), Kobe, Hyogo 650-0047, Japan
${}^{4}$ Interdisciplinary Theoretical and Mathematical Sciences Program (iTHEMS), Wako, Saitama 351-0198, Japan
${}^{5}$omputational Condensed Matter Physics Laboratory, RIKEN Cluster for Pioneering Research (CPR), Saitama 351-0198, Japan
@@
To assess the accuracy of various tensor network methods, including matrix product state (MPS) [2] and two-dimensional tensor network state (2dTNS) methods (for details of our 2dTNS simulations, see Sec. S5), we compare the results of $\langle\hat{Z}_{\rm avg}(t)\rangle$ obtained by these methods with the numerically exact results calculated by the state-vector method in Figs. S4 and S5. As shown in Figs. S4a-c and S5a-c, when $\theta_{x}=0.9\pi$, $\langle\hat{Z}_{\rm avg}(t)\rangle$ obtained by these tensor network methods with relatively small bond dimensions $\chi$ (MPS with $\chi=300$ and 2dTNS with $\chi=40$) sufficiently converge consistently to the numerically exact values, even for time steps up to 100. However, as shown in Figs. S4d-f and S5d-f, when $\theta_{x}=0.8\pi$, which is close to the crossover boundary between the DTC/DTQC and thermalised regimes, a larger $\chi$ is required to obtain the converged value, due to the increase of entanglement. For example, when $(\theta_{x},\theta_{z})=(0.8\pi,0)$ (Fig. S5d), 2dTNS with $\chi=400$ is insufficient to obtain the numerically exact results for $t\gtrsim 50$. A similar trend is observed for other parameters shown in Figs. S5e,f and for the MPS method shown in Figs. S4d-f. These results suggest that the long-time time-evolution simulations of $\langle\hat{Z}_{\rm avg}(t)\rangle$ based on classical tensor network methods already face challenges even for $L=28$ in a parameter region around $\theta_{x}=0.8\pi$. We expect to encounter similar challenges for $L=133$, where the numerically exact state-vector method cannot be applied, although a recent report suggests that the accuracy of a 2dTNS method improves as the system size increases [3]. This emphasises the significance of utilising a quantum computer precisely in this parameter regime.
@@
Figure S6 shows the error-unmitigated raw data of the average magnetisation $\langle\hat{Z}_{\rm avg}(t)\rangle_{0}$ for up to 100 time steps obtained for the $L=133$ system using ibm_torino (see Fig. 1a). Here, the average is performed over the same set $A$ of qubits as described in the main text, i.e., $A=\{57,58,59,60,61,62,63,64,65,66,67,68,69,70,71\}$. Remarkably, even without error mitigation, distinct signals of period-doubling oscillations persist for over 100 time steps. To mitigate errors in the raw data for $\theta_{x}\neq\pi$, the results corresponding to the trivial cases with $\theta_{x}=\pi$ are utilised, as explained in Eq. (4).
@@
Figure S6: Error-unmitigated raw data $\langle\hat{Z}_{\rm avg}(t)\rangle_{0}$ for the heavy-hexagonal lattice of $L=133$ qubits obtained using ibm_torino (see Fig. 1a). The parameters $(\theta_{x},\theta_{z})$ are a $(\pi,0)$, b $(\pi,0.25\pi)$, c $(\pi,0.5\pi)$, d $(\pi,0.75\pi)$, e $(\pi,\pi)$, f $(0.9\pi,0)$, g $(0.9\pi,0.25\pi)$, h $(0.9\pi,0.5\pi)$, i $(0.9\pi,0.75\pi)$, j $(0.9\pi,\pi)$, k $(0.85\pi,0)$, l $(0.85\pi,0.25\pi)$, m $(0.85\pi,0.5\pi)$, n $(0.85\pi,0.75\pi)$, o $(0.85\pi,\pi)$, p $(0.8\pi,0)$, q $(0.8\pi,0.25\pi)$, r $(0.8\pi,0.5\pi)$, s $(0.8\pi,0.75\pi)$, and t $(0.8\pi,\pi)$.
@@
For the trivial cases with $\theta_{x}=\pi$, as shown in Figs. S6a-e, $|\langle\hat{Z}_{\rm avg}(t)\rangle_{0}|$ should ideally be 1. However, deviations from this ideal value are observed, which become more pronounced with increasing time steps, owing to the noise and decoherence inherent in the quantum device. The number of CZ gates in the quantum circuit increases by 150 for each operation of $\hat{U}_{\text{F}}$ as defined in Eq. (2). At $t/T=100$, the quantum circuit volume $v=15000$ (defined by the total number of CZ gates) superficially implies $|\langle\hat{Z}_{\rm avg}(t=100T)\rangle_{0}|=(1-p)^{v}\simeq 10^{-27}$, with a typical two-qubit infidelity in ibm_torino of $p=4\times 10^{-3}$. However, in practice, we observe $|\langle\hat{Z}_{\rm avg}(t=100T)\rangle_{0}|\simeq 0.1$ in Figs. S6a-e, suggesting an effective quantum circuit volume $v_{\text{eff}}\simeq 600$. Since we measure local observables, $v_{\text{eff}}$ is significantly smaller than $v$, as discussed in Ref. [4]. Additionally, it is interesting to notice that the raw data for the $L=28$ system with the trivial parameters, as shown in Figs. S1a-c, also exhibit $|\langle\hat{Z}_{\rm avg}(t=100T)\rangle_{0}|\simeq 0.1-0.15$, indicating a similar $v_{\text{eff}}$.
@@
In Fig. 2d of the main text, we demonstrated that the averaged magnetisation $\langle\hat{Z}_{\rm avg}(t)\rangle$ obtained from ibm_torino with the error mitigation agrees well with those of classical 2dTNS simulations at $(\theta_{x},\theta_{z})=(0.8\pi,0.5\pi)$ for the $L=133$ system over 50 time steps. In Figs. S7 and S8, we extend this comparison to other parameters over 100 time steps. As shown in Figs. S7a-c and S8a-c, when $\theta_{x}=0.9\pi$, the averaged magnetisation $\langle\hat{Z}_{\rm avg}(t)\rangle$ obtained by the 2dTNS method with a relatively small bond dimension $\chi=20$ already converges sufficiently. This implies that the 2dTNS results for these parameter sets are reliable even for up to 100 time steps. Indeed, this parameter region exhibits limited entanglement growth over time steps, as indicated by the agreement between the MPS results with relatively small bond dimensions and the 2dTNS results (see Figs. S9a-c and S10a-c). Moreover, it is particularly remarkable that despite the simple error mitigation protocol introduced in Eq. (4) of the main text, the error-mitigated $\langle\hat{Z}_{\rm avg}(t)\rangle$ obtained from ibm_torino are in relatively good alignment with the 2dTNS results (see Figs. S7a-c and S8a-c).
@@
On the other hand, as shown in Figs. S7d-f and S8d-f, the convergence of the 2dTNS results for $\theta_{x}=0.8\pi$ is slow with increasing the bond dimension, especially for long time steps, as anticipated from the results for the $L=28$ system shown in Figs. S5d-f. This parameter region is close to the crossover boundary between the DTC/DTCQ and thermalised regimes, where extensive entanglement growth is expected. Due to the substantial increase in entanglement, the performance of tensor-network simulations is significantly degraded, resulting in a noticeable discrepancy between MPS and 2dTNS as shown in Figs. S9d-f and S10d-f. Although the classical 2dTNS simulations may not have yet converged in these parameter sets for $t\gtrsim 50$, we still observe that the error-mitigated $\langle\hat{Z}_{\rm avg}(t)\rangle$ obtained from ibm_torino align well with the 2dTNS results (see Figs. S7d-f and S8d-f). Given that this is the parameter region where classical simulations face challenges, it underscores the value of quantum computers precisely in this parameter region.
@@
Figure S7: Error-mitigated data $\langle\hat{Z}_{\rm avg}(t)\rangle$ are compared with the results obtained by the 2dTNS method with various bond dimensions ($\chi=10$, 20, 100, 200, and 300) for the heavy-hexagonal lattice of $L=133$ qubits. The parameters $(\theta_{x},\theta_{z})$ are a $(0.9\pi,0)$, b $(0.9\pi,0.5\pi)$, c $(0.9\pi,\pi)$, d $(0.8\pi,0)$, e $(0.8\pi,0.5\pi)$, and f $(0.8\pi,\pi)$.
@@
Figure S9: Comparison of the results for $\langle\hat{Z}_{\rm avg}(t)\rangle$ on the heavy-hexagonal lattice of $L=133$ qubits obtained by the 2dTNS method with a given bond dimension ($\chi=20$ in panels a-c and 300 in panels d-f) and the MPS method with various bond dimensions ($\chi=1000$, 3000, and 5000). The parameters $(\theta_{x},\theta_{z})$ are a $(0.9\pi,0)$, b $(0.9\pi,0.5\pi)$, c $(0.9\pi,\pi)$, d $(0.8\pi,0)$, e $(0.8\pi,0.5\pi)$, and f $(0.8\pi,\pi)$.

### 12 Mar 2026 v2

@@ Abstract
Figure 3: Dynamics of magnetisation exhibiting DTC, IM-DTC, and thermalisation. The raw data $\langle\hat{Z}_{\rm avg}(t)\rangle_{0}$ (cyan circles) and the error-mitigated data $\langle\hat{Z}_{\rm avg}(t)\rangle$ (yellow diamonds) obtained on the heavy-hexagonal lattice of $L=133$ qubits at various parameter sets $(\theta_{x},\theta_{z})$: a $(0.9\pi,0)$, b $(0.9\pi,0.5\pi)$, c $(0.9\pi,\pi)$, d $(0.85\pi,0)$, e $(0.85\pi,0.5\pi)$, f $(0.85\pi,\pi)$, g $(0.8\pi,0)$, h $(0.8\pi,0.5\pi)$, i $(0.8\pi,\pi)$, j $(0.7\pi,0.0\pi)$, k $(0.7\pi,0.5\pi)$, and l $(0.7\pi,\pi)$. Notice that the time duration in the horizontal axis in j-l is half of that in the other panels.
The clean DTC phase retains its stability even with the addition of longitudinal field $\theta_{z}$, which explicitly breaks Ising symmetry. Here, the DTC order results from the spontaneous breaking of an emergent Ising symmetry [13, 16]. In addition to the period-doubling oscillation, $\theta_{z}$ induces a longer-period oscillation, as clearly seen in Fig. 3 (also see Fig. S6).
To analyse this additional modulation, we perform a discrete Fourier transform of the error-mitigated magnetisation, $\tilde{Z}(\omega)=\bigl|\frac{1}{n_{\mathrm{max}}}\sum_{n=0}^{n_{\mathrm{max}}-1}\langle\hat{Z}_{\rm avg}(nT)\rangle e^{-i\omega nT}\bigr|$, where $\omega T=2\pi k/n_{\mathrm{max}}$ with $k=0,1,\ldots,n_{\mathrm{max}}-1$, and $n_{\mathrm{max}}=100$. As shown in Figs. 4a–4c, when $\theta_{z}=0$, the Fourier spectrum exhibits only a single peak at $\omega T/(2\pi)=0.5$, consistent with a pure period-doubling DTC response. Introducing a longitudinal field produces additional side peaks, symmetrically located at frequencies $\omega_{\pm}$ around the DTC peak. These side peaks shift systematically with $\theta_{z}$, reflecting a continuous and tunable detuning of the modulation frequency.
Although the discrete Fourier frequencies $\omega_{\pm}T/(2\pi)$ are rational numbers by definition, their smooth dependence on $\theta_{z}$ indicates that, for fixed parameters and increasing observation time, the corresponding frequencies fluctuate around generally irrational values. The appearance of these additional, parameter-sensitive and essentially incommensurate frequency components constitutes an IM-DTC response, reminiscent of the behaviour analysed in Refs. [24, 27].
@@
Figure 4: Fourier analysis of incommensurately modulated DTC responses induced by the longitudinal field. Fourier spectrum $\tilde{Z}(\omega)$ of the averaged magnetisation for a $\theta_{x}=0.9\pi$, b $\theta_{x}=0.85\pi$, and c $\theta_{x}=0.8\pi$ with various values of $\theta_{z}=0,0.125\pi,0.25\pi,\cdots,\pi$ (from bottom to top). The spectra are vertically offset by 0.05 for clarity. d, Envelope frequency $\omega_{\text{env}}$ of the IM-DTC modulation as a function of the perturbation parameter $\epsilon$ to the transverse field. e, Intensity of the side peaks, $A_{\text{side}}$, as a function of $\theta_{z}$. A schematic diagram illustrating the crossover between a clean DTC response and an IM-DTC response is shown at the top of the figure. Error bars in a-c and e represent the propagated error due to sampling errors in $\tilde{Z}(\omega)$. However, they are smaller than the size of symbols.
@@
The envelope frequency of the longer-period IM-DTC oscillation can be estimated as $\omega_{\text{env}}=(\omega_{+}-\omega_{-})/2$. We observe that the envelope frequency increases proportionally to the perturbation to the transverse field $h_{x}$, i.e., $\omega_{\text{env}}\propto\epsilon$ (Fig. 4d). Particularly at $\theta_{z}=\pi$, $\omega_{\text{env}}T/(2\pi)=0.05$, $0.07$, and $0.09$ for $\theta_{x}=0.9\pi$, $0.85\pi$, and $0.8\pi$, respectively. These values approximately follow $\omega_{\text{env}}T/(2\pi)\simeq\epsilon/(2\theta_{J})$, which is proportional to the frequency of a Bloch oscillation [51] induced by a weak transverse field in a one-dimensional Ising model with long-range interactions [52]. In contrast to the frequency of a period-doubling oscillation, the incommensurate frequency of a side peak $\omega_{\pm}$ is sensitive to $\epsilon$. This IM-DTC behavior with modulated period-doubling oscillations induced by a single-color periodic driving leads to the same type of IM-DTC response as proposed in Refs. [24, 27], but differs from the one in Refs. [53, 54, 55].
@@
To characterise the crossover between a DTC and an IM-DTC response, we plot the sum of the intensities $A_{\pm}$ of the side peaks at $\omega_{\pm}$, denoted as $A_{\text{side}}=A_{+}+A_{-}$, in Fig. 4e. $A_{\mathrm{side}}$ tends to decrease with decreasing $\theta_{x}$. At $\theta_{x}=0.7\pi$, $\langle\hat{Z}_{\rm avg}(t)\rangle$ quickly decays with increasing time step $t/T$ as seen in Figs. 3j-3l, indicating the absence of both a long-lived DTC and an IM-DTC modulation. When $\theta_{x}=\pi$, only the period-doubling DTC is present, as $|\langle\hat{Z}_{\rm avg}(t)\rangle|$ remains 1 regardless of the value of $\theta_{z}$. Therefore, we conclude that an IM-DTC response manifests in a prethermal regime persisting within the timescale $0\leq t/T\leq 100$ for a parameter range of $0.8\pi\lesssim\theta_{x}\lesssim 0.9\pi$ and $0.25\pi\lesssim\theta_{z}\leq\pi$.
@@
The recent advancements in the quality of quantum devices have enabled the successful simulation of quantum dynamics on a much larger scale, both in terms of qubit count and time step duration. Specifically, achieving quantum dynamics simulation on a 133-qubit system for up to 100 time steps using a digital quantum computer, as demonstrated in this study, represents a significant leap forward compared to previous studies. Our comparative study, employing a 2dTNS method, confirmed the agreement between the results obtained using the quantum device and those from the classical simulations, for approximately up to 50 time steps. However, in our investigation of both 28-qubit and 133-qubit systems, we encountered a critical parameter regime where the perturbation to the transverse field approaches $2\epsilon\approx 0.2\pi$, near the crossover boundary between DTC and IM-DTC responses and thermalisation. In this regime, the limitation of classical simulations becomes apparent, particularly for long-time steps around 100. Indeed, in the 28-qubit system, we observe discrepancies between the results of the 2dTNS method, even with the largest feasible bond dimension, and those of the state-vector simulation (Supplementary Information S2). Additionally, in the 133-qubit system, a slow convergence of the results of the 2dTNS method with respect to the bond dimension is observed, implying the growth of entanglement that is hardly tractable within the classical resources available for the 2dTNS method (Supplementary Information S3). Therefore, our study of Floquet dynamics in this parameter regime, extending up to 100 time steps on the 133-qubit system, pushes the boundaries of classical simulations specially based on tensor-network methods to their limit, emphasising the significant potential of current digital quantum computers for simulating out-of-equilibrium quantum dynamics in two dimensions. The utilisation of a stabilizer formalism may further improve the simulatability of a tensor network [56, 57]. It would be a very interesting future challenge to reproduce the DTC and IM-DTC dynamics in the prethermal regime up to 100 time steps using a tensor network method that incorporates quantum noise [58].
@@
Supplementary Information to “Unveiling clean two-dimensional discrete time crystals on a digital quantum computer”
Kazuya Shinjo1, Kazuhiro Seki2, Tomonori Shirakawa2,3,4,5, Rong-Yang Sun2,3,4, and Seiji Yunoki1,2,3,5
1Computational Quantum Matter Research Team, RIKEN Center for Emergent Matter Science (CEMS), Wako, Saitama 351-0198, Japan
2Quantum Computational Science Research Team, RIKEN Center for Quantum Computing (RQC), Wako, Saitama 351-0198, Japan
3Computational Materials Science Research Team, RIKEN Center for Computational Science (R-CCS), Kobe, Hyogo 650-0047, Japan
4RIKEN Interdisciplinary Theoretical and Mathematical Sciences Program (iTHEMS), Wako, Saitama 351-0198, Japan
5Computational Condensed Matter Physics Laboratory, RIKEN Cluster for Pioneering Research (CPR), Saitama 351-0198, Japan
@@
To assess the accuracy of various tensor network methods, including matrix product state (MPS) [2] and two-dimensional tensor network state (2dTNS) methods (for details of our 2dTNS simulations, see Sec. S9), we compare the results of $\langle\hat{Z}_{\rm avg}(t)\rangle$ obtained by these methods with the numerically exact results calculated by the state-vector method in Figs. S4 and S5. As shown in Figs. S4a-S4c and S5a-S5c, when $\theta_{x}=0.9\pi$, $\langle\hat{Z}_{\rm avg}(t)\rangle$ obtained by these tensor network methods with relatively small bond dimensions $\chi$ (MPS with $\chi=300$ and 2dTNS with $\chi=40$) sufficiently converge consistently to the numerically exact values, even for time steps up to 100. However, as shown in Figs. S4d-S4f and S5d-S5f, when $\theta_{x}=0.8\pi$, which is close to the crossover boundary between the discrete time crystal (DTC) and incommensurately modulated DTC (IM-DTC) responses and the thermalised regime, a larger $\chi$ is required to obtain the converged value, due to the increase of entanglement. For example, when $(\theta_{x},\theta_{z})=(0.8\pi,0)$ (Fig. S5d), 2dTNS with $\chi=400$ is insufficient to obtain the numerically exact results for $t\gtrsim 50$. A similar trend is observed for other parameters shown in Figs. S5e and S5f and for the MPS method shown in Figs. S4d-S4f. These results suggest that the long-time time-evolution simulations of $\langle\hat{Z}_{\rm avg}(t)\rangle$ based on classical tensor network methods already face challenges even for $L=28$ in a parameter region around $\theta_{x}=0.8\pi$. We expect to encounter similar challenges for $L=133$, where the numerically exact state-vector method cannot be applied, although a recent report suggests that the accuracy of a 2dTNS method improves as the system size increases [3]. This emphasises the significance of utilising a quantum computer precisely in this parameter regime.
@@
Figure S6 shows the error-unmitigated raw data of the average magnetisation $\langle\hat{Z}_{\rm avg}(t)\rangle_{0}$ and the error-mitigated data $\langle\hat{Z}_{\rm avg}(t)\rangle$ for up to 100 time steps obtained for the $L=133$ system using ibm_torino (see Fig. 1a). Here, the average is performed over the same set $A$ of qubits as described in the main text, i.e., $A=\{57,58,59,60,61,62,63,64,65,66,67,68,69,70,71\}$. Remarkably, even without error mitigation, distinct signals of period-doubling oscillations persist for over 100 time steps. To mitigate errors in the raw data for $\theta_{x}\neq\pi$, the results corresponding to the trivial cases with $\theta_{x}=\pi$ are utilised, as explained in Eq. (4).
@@
Figure S6: Error-unmitigated raw data $\langle\hat{Z}_{\rm avg}(t)\rangle_{0}$ (cyan circles) and error-mitigated data $\langle\hat{Z}_{\rm avg}(t)\rangle$ (yellow diamonds) for the heavy-hexagonal lattice of $L=133$ qubits obtained using ibm_torino (see Fig. 1a). The parameters $(\theta_{x},\theta_{z})$ are a $(\pi,0)$, b $(\pi,0.25\pi)$, c $(\pi,0.5\pi)$, d $(\pi,0.75\pi)$, e $(\pi,\pi)$, f $(0.9\pi,0)$, g $(0.9\pi,0.25\pi)$, h $(0.9\pi,0.5\pi)$, i $(0.9\pi,0.75\pi)$, j $(0.9\pi,\pi)$, k $(0.85\pi,0)$, l $(0.85\pi,0.25\pi)$, m $(0.85\pi,0.5\pi)$, n $(0.85\pi,0.75\pi)$, o $(0.85\pi,\pi)$, p $(0.8\pi,0)$, q $(0.8\pi,0.25\pi)$, r $(0.8\pi,0.5\pi)$, s $(0.8\pi,0.75\pi)$, and t $(0.8\pi,\pi)$. The error-mitigated data for the cases of $\theta_{x}=\pi$ are not shown because they are $\langle\hat{Z}_{\rm avg}(t)\rangle=\pm 1$ by definition.
@@
For the trivial cases with $\theta_{x}=\pi$, as shown in Figs. S6a-S6e, $|\langle\hat{Z}_{\rm avg}(t)\rangle_{0}|$ should ideally be 1. However, deviations from this ideal value are observed, which become more pronounced with increasing time steps, owing to the noise and decoherence inherent in the quantum device. The number of CZ gates in the quantum circuit increases by 150 for each operation of $\hat{U}_{\text{F}}$ as defined in Eq. (2). At $t/T=100$, the quantum circuit volume $v=15000$ (defined by the total number of CZ gates) superficially implies $|\langle\hat{Z}_{\rm avg}(t=100T)\rangle_{0}|=(1-p)^{v}\simeq 10^{-27}$, with a typical two-qubit infidelity in ibm_torino of $p=4\times 10^{-3}$. However, in practice, we observe $|\langle\hat{Z}_{\rm avg}(t=100T)\rangle_{0}|\simeq 0.1$ in Figs. S6a-S6e, suggesting an effective quantum circuit volume $v_{\text{eff}}\simeq 600$. Since we measure local observables, $v_{\text{eff}}$ is significantly smaller than $v$, as discussed in Ref. [4]. Additionally, it is interesting to notice that the raw data for the $L=28$ system with the trivial parameters, as shown in Figs. S1a-S1c, also exhibit $|\langle\hat{Z}_{\rm avg}(t=100T)\rangle_{0}|\simeq 0.1-0.15$, indicating a similar $v_{\text{eff}}$.
@@
In Fig. 2d of the main text, we demonstrated that the averaged magnetisation $\langle\hat{Z}_{\rm avg}(t)\rangle$ obtained from ibm_torino with the error mitigation agrees well with those of classical 2dTNS simulations at $(\theta_{x},\theta_{z})=(0.8\pi,0.5\pi)$ for the $L=133$ system over 50 time steps. In Figs. S7 and S8, we extend this comparison to other parameters over 100 time steps. As shown in Figs. S7a-S7c and S8a-S8c, when $\theta_{x}=0.9\pi$, the averaged magnetisation $\langle\hat{Z}_{\rm avg}(t)\rangle$ obtained by the 2dTNS method with a relatively small bond dimension $\chi=20$ already converges sufficiently. This implies that the 2dTNS results for these parameter sets are reliable even for up to 100 time steps. Indeed, this parameter region exhibits limited entanglement growth over time steps, as indicated by the agreement between the MPS results with relatively small bond dimensions and the 2dTNS results (see Figs. S9a-S9c and S10a-S10c). Moreover, it is particularly remarkable that despite the simple error mitigation protocol introduced in Eq. (4) of the main text, the error-mitigated $\langle\hat{Z}_{\rm avg}(t)\rangle$ obtained from ibm_torino are in relatively good alignment with the 2dTNS results (see Figs. S7a-S7c and S8a-S8c).
@@
On the other hand, as shown in Figs. S7d-S7f and S8d-S8f, the convergence of the 2dTNS results for $\theta_{x}=0.8\pi$ is slow with increasing the bond dimension, especially for long time steps, as anticipated from the results for the $L=28$ system shown in Figs. S5d-S5df. This parameter region is close to the crossover boundary between the DTC and IM-DTC responses and the thermalised regime, where extensive entanglement growth is expected. Due to the substantial increase in entanglement, the performance of tensor-network simulations is significantly degraded, resulting in a noticeable discrepancy between MPS and 2dTNS as shown in Figs. S9d-S9f and S10d-S10f. Although the classical 2dTNS simulations may not have yet converged in these parameter sets for $t\gtrsim 50$, we still observe that the error-mitigated $\langle\hat{Z}_{\rm avg}(t)\rangle$ obtained from ibm_torino align well with the 2dTNS results (see Figs. S7d-S7f and S8d-S8f). Given that this is the parameter region where classical simulations face challenges, it underscores the value of quantum computers precisely in this parameter region.
@@
Figure S7: Error-mitigated data $\langle\hat{Z}_{\rm avg}(t)\rangle$ are compared with the results obtained by the 2dTNS method with various bond dimensions ($\chi=10$, 20, 100, 200, 300, and 400) for the heavy-hexagonal lattice of $L=133$ qubits. The parameters $(\theta_{x},\theta_{z})$ are a $(0.9\pi,0)$, b $(0.9\pi,0.5\pi)$, c $(0.9\pi,\pi)$, d $(0.8\pi,0)$, e $(0.8\pi,0.5\pi)$, and f $(0.8\pi,\pi)$.
@@
Figure S9: Comparison of the results for $\langle\hat{Z}_{\rm avg}(t)\rangle$ on the heavy-hexagonal lattice of $L=133$ qubits obtained by the 2dTNS method with a given bond dimension ($\chi=20$ in panels a-c and 400 in panels d-f) and the MPS method with various bond dimensions ($\chi=1000$, 3000, and 5000). The parameters $(\theta_{x},\theta_{z})$ are a $(0.9\pi,0)$, b $(0.9\pi,0.5\pi)$, c $(0.9\pi,\pi)$, d $(0.8\pi,0)$, e $(0.8\pi,0.5\pi)$, and f $(0.8\pi,\pi)$.